Ocean Fishers Ltd had a 22-foot fishing boat with an inboard motor that was purchased on April 9, 2009, for $86,500. The PPE Sub ledger shows the following information regarding the boat: Fishing Boat-22 Foot With Inboard Motor Component Fibreglass body Date of Purchase Apr. 9/09 Apr. 9/09 Deprec. Method SL SL Cost s 25.500 S 6,000 Est. Residual Est. Life 15 yr Motor 61,000 10 yr S 86,500 On June 27, 2017, $67,000 cash was paid for a new motor to replace the old one, which was scrapped. The new motor had an estimated useful life of 12 years and a residual value of $4,000. Early in 2017, it was determined that the useful life of the boats fibreglass body should be adjusted to a total of 20 years with no change in the residual value Required 1. Record the appropriate entries regarding the a. Purchase of the replacement motor on June 27, 2017. (If no entry is required for a transaction, select No journal entry required in the first account field.)

b. Depreciation taken on the fishing boat (body plus motor) on December 31, 2017, the companys year-end. (If no entry is required for a transaction, select No journal entry required in the first account field. Round the final answers to the nearest whole dollar.)

2. Calculate total depreciation taken on the fishing boat (body plus motor) for the companys year ended December 31, 2017. (Round the final answers to the nearest whole dollar.) Total 2017 depreciation
Answer #1

Answer:

Component Fibreglss Old Motor New Motor
Purchase date 09-Apr-09 09-04-2009 27-Jun-17
Cost 25,500 61,000 67,000
Estimated Residual 6,000 4,000
Estimated life 15 10 12
Depreciation=(Cost - Salvage value) Est.life
Annual Depreciation
Fibreglss(25500-6000)15                                                        1,100
Old Motor(61,000/10)                                                        6,100
New Motor(67000-4,000)/12                                                        5,250
Depreciation
2009                                                            825                   4,575 =6,100/12*9 for 9 months from 9 April 2009
2010-2016                               7,700                 42,700 =6100*7 for 7 years
2017till 27June 2017                                   3,050 =6,100/12*6 for 6 months ti ll 27 June 2017
Accumulated Dep                               8,525                 50,325
Net Fixed Asset                             16,975                 10,675 =61,000-50,325
Scrapped                                          -  
Loss on scrapped motor                                 10,675
Est. life - revised 20
Est. remaining life -Old 7.25 =15 years -7.75 Years
Est. remaining life - revised 12.25 =7.25 years -5 Years
Depreciation revised =(Net Fixed Asset - Salvage value) Est.revised life
=(16,975-6000)/12.25
Depreciation Fibreglss Old Motor
2017                                                      895.91                                   2,650 =5250/12*6
Accumulated Dep                                                  9,420.91                                   2,650
Journal entries
Date Account Debit($) Credit($)
27-Jun-17 Depreciation expense - Old Motor                 3,050
                    Accumulated Depreciation                                    3,050
( To record depreciation in 2017 regarding motor being replaced)
27-Jun-17 New Motor               67,000
Accumulated Depreciation               50,325
Loss on scrapped asset               10,675
             Old Motor                                  61,000
            Cash                                  67,000
( To record purchase of the replacement motor and old motor scrapped)
31-Dec-17 Depreciation expense - Boat body                     896
Depreciation expense - New Motor                 2,650
Accumulated Depreciation                                    3,546
( To record revised depreciation expense on fishing boat)
31-Dec-17 Depreciation expense - Boat body                     896
Depreciation expense - New Motor                 2,650
Depreciation expense - Old Motor                 3,050                                    6,596
               Accumulated Depreciation
(To record total depreciation on thefishing boat)